OpenStudy (anonymous):

The following is a geometric sequence. -1/2, 1/4, -1/8, 1/16

OpenStudy (anonymous):

@Loser66

OpenStudy (anonymous):

idk im confuessed

OpenStudy (michele_laino):

hint: the ratio between one term and its preceding term is constant, and we can write this: \[\left( {1/16} \right):\left( { - 1/8} \right) = \left( { - 1/8} \right):\left( {1/4} \right) = \left( {1/4} \right):\left( { - 1/2} \right) = ...?\]

OpenStudy (anonymous):

still lost.

OpenStudy (michele_laino):

hint: \[\frac{1}{{16}}:\left( { - \frac{1}{8}} \right) = \frac{1}{{16}} \times \left( { - 8} \right) = ...\]

OpenStudy (anonymous):

i do.. and @Michele_Laino is it -1/2?

OpenStudy (michele_laino):

yes! that's right!

OpenStudy (anonymous):

so i mutliply each by -1/2?

OpenStudy (michele_laino):

yes you have to multiply one term by -1/2 in order to get the subsequent term of your geometric sequence

OpenStudy (anonymous):

so its only a geometric sequence if its multiply by a fraction?

OpenStudy (michele_laino):

no, the constant can be also an integer number, or even a irrational number

OpenStudy (anonymous):

so is something like -5,0,5,10 a geometric sequence?

OpenStudy (michele_laino):

no, it is an arithmetic sequence, since the difference between one term and its preceding term is constant: 0-(-5)= 5-0= 10-5=...?

OpenStudy (anonymous):

okay! but what if its like -5,25,-125,625 when its all multiply by -5

OpenStudy (michele_laino):

it is a geometric sequence whose constant is -5

OpenStudy (anonymous):

but why?

OpenStudy (michele_laino):

because the ration between one term and its preceding term is a constant value

OpenStudy (michele_laino):

ratio*

OpenStudy (anonymous):

okay i get it

OpenStudy (michele_laino):

ok!